A car travels at 15 m/s for 10 s. It then speeds up with a constant acceleration of 2.0 m/s 2 for 15 s. At the end of this time, what is its velocity?


1
Expert's answer
2021-02-02T09:30:37-0500
"v_f=v_i+at\\v_f=15+2(15)=45\frac{m}{s}"